OpenStudy (anonymous):

Arrange the steps to solve this system of linear equations in the correct sequence. x + y = -2 2x – 3y = -9

OpenStudy (anonymous):

you solve it by doing substitution

OpenStudy (anonymous):

here's an example to help x+y=5 y=x+5 x-2y=2 x+2(x+5)=2 x+2x+10=2 3x+10=2 3x=-12 X=-4